# Monte-Carlo Integration Based Multiple-Scattering Channel Modeling for Ultraviolet Communications in Turbulent Atmosphere

### v1

@@ IV-C Estimating the Turbulent Fluctuation
where we have used the fact that all sampling points $\{\bm{s}_{n}^{1},\bm{s}_{n}^{2},\cdots,\bm{s}_{n}^{M}\}$ are independent from each other and $\text{Var}[O_{n}(\bm{s}_{n}^{1})]=\text{Var}[O_{n}(\bm{s}_{n}^{2})]\cdots=\text{Var}[O_{n}(\bm{s}_{n}^{M})]=\text{Var}[O_{n}(\bm{d},\bm{\theta},\bm{\phi})]$ is the variance of the objective function. We can see that the variance $\text{Var}[P_{tur,n}]$ depends on both $\text{Var}[O_{n}(\bm{d},\bm{\theta},\bm{\phi})]$ and $M$. Besides, we have $\text{Var}[P_{tur,n}]\to 0$ when the number of sampling points $M\to\infty$, which means the calculated variance can be arbitrarily small as long as $M$ is large enough. Whereas, larger sampling points $M$ in the Monte-Carlo process only reduce the fluctuation caused by randomness, leading to more stable and precise results. Therefore, any attempt to calculate the turbulent fluctuation by estimating $\text{Var}[P_{tur,n}]$ will fail due to the fact that $\text{Var}[P_{tur,n}]$ depends on $M$. Here we present an improved estimating method for the turbulent fluctuation, which is independent of $M$.
By substituting (16) into (18), we can obtain
$P_{tur,n}(\bm{\eta},\bm{d},\bm{\theta},\bm{\phi})=\begin{cases}\prod_{i=0}^{n}\eta_{i}P_{d,n}(\bm{d},\bm{\theta},\bm{\phi}),\quad\{\bm{d},\bm{\theta},\bm{\phi}\}\in\mathcal{F},\\ 0,\quad\{\bm{d},\bm{\theta},\bm{\phi}\}\notin\mathcal{F},\end{cases}$ (38)
@@
$P_{tur,n}(\bm{\eta})$ $=\!\prod_{i=0}^{n}\eta_{i}\!\!\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}\!P_{n}(\bm{d},\bm{\theta},\bm{\phi})f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}=\!{\prod_{i=0}^{n}\eta_{i}}P_{n}.$ (39)
Then we can define an equivalent turbulent fading coefficient $\eta_{eq,n}\triangleq P_{tur,n}(\bm{\eta})/P_{n}$ on the average receiving power $P_{n}$ such that the instantaneous receiving power $P_{tur,n}(\bm{\eta})=\eta_{eq,n}P_{n}$. Obviously, we have
@@
$\sigma_{tur,n}^{2}$ $\triangleq\text{Var}[\eta_{eq,n}]=\int_{\bm{\eta}}(\eta_{eq,n}-1)^{2}f_{H}(\bm{\eta})\mathrm{d}\bm{\eta},$ (41)
where $f_{H}(\bm{\eta})$ is the joint PDF of $\{\eta_{0},\eta_{1},\cdots,\eta_{n}\}$. From (38) we can see that those propagating paths with $\{\bm{d},\bm{\theta},\bm{\phi}\}\notin\mathcal{F}$ contribute zero to the turbulent fluctuation effect. Therefore, we should restrict the variance on the set $\mathcal{F}$. Then the joint PDF $f_{H}(\bm{\eta})$ can be obtained by averaging $\{\bm{d},\bm{\theta},\bm{\phi}\}$ out of the conditional PDF $f_{H|D,\Theta,\Phi}(\bm{\eta}|\bm{d},\bm{\theta},\bm{\phi})$ in (20) on the set $\mathcal{F}$, i.e.,
$f_{H}(\bm{\eta})$ $\!=\!\int_{\mathcal{F}}f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi}|\{\bm{d},\bm{\theta},\bm{\phi}\}\in\mathcal{F})f_{H|D,\Theta,\Phi}(\bm{\eta}|\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi},$ (42)
where $f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi}|\{\bm{d},\bm{\theta},\bm{\phi}\}\in\mathcal{F})$ is the joint PDF of $\{\bm{d},\bm{\theta},\bm{\phi}\}$ on the set $\mathcal{F}$; and we have
$f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi}|\{\bm{d},\bm{\theta},\bm{\phi}\}\in\mathcal{F})$ $=\frac{f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})}{\text{Pr}[\{\bm{d},\bm{\theta},\bm{\phi}\}\in\mathcal{F}]},$ (43)
@@
$\text{Pr}[\{\bm{d},\bm{\theta},\bm{\phi}\}\in\mathcal{F}]=\int_{\mathcal{F}}f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}.$ (44)
@@
$f_{H}(\bm{\eta})$ $=\frac{\int_{\mathcal{F}}\prod_{i=0}^{n}f_{H}(\eta_{i};d_{i})f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}}{\int_{\mathcal{F}}f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}}$ (45)
$=\frac{\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}\prod_{i=0}^{n}f_{H}(\eta_{i};d_{i})I_{n}f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}}{\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}I_{n}f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}}.$
@@
$\sigma_{tur,n}^{2}$ $\!=\!\frac{\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}(\prod_{i=0}^{n}M_{2}(d_{i})-1)I_{n}f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}}{\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}I_{n}(\bm{d},\bm{\theta},\bm{\phi})f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}},$ (46)
@@
$\sigma_{tur,n}^{2}$ $=\frac{\frac{1}{M}\sum_{m=1}^{M}\left[\prod_{i=0}^{n}M_{2}(d_{i}^{m})-1\right]I_{n}(\bm{s}_{n}^{m})}{\frac{1}{M}\sum_{m=0}^{M}I_{n}(\bm{s}_{n}^{m})}=\frac{1}{Count}\sum_{m=1}^{M}\left[\prod_{i=0}^{n}M_{2}(d_{i}^{m})-1\right]I_{n}(\bm{s}_{n}^{m}),$ (47)
@@
Similar to the $n$th scattering order, the total turbulent fluctuation effect on the total receiving power of $N$ scattering orders can be characterized by defining an equivalent turbulent fading coefficient $\eta_{eq}\triangleq\sum_{n=1}^{N}P_{tur,n}(\bm{\eta})/\sum_{n=1}^{N}P_{n}$. Using the relations $P_{tur,n}(\bm{\eta})=\eta_{eq,n}P_{n}$ and $P_{tot}=\sum_{n=1}^{N}P_{n}$, we can obtain
@@
IV-D Estimating the Distribution of Turbulent Fading Coefficient
It is challenging to derive the distribution of the turbulent fading coefficient for a general turbulent fading model. However, when the turbulence is weak and an LN fading is considered, then we can further derive the distributions of the equivalent turbulent fading coefficients $\eta_{eq,n}$ and $\eta_{eq}$. For a given propagating path $\{\bm{d},\bm{\theta},\bm{\phi}\}$, the equivalent turbulent fading coefficient $\eta_{eq,n}$ can be obtained from (40) as
@@ IV-D Estimating the Distribution of Turbulent Fading Coefficient
Because $\eta_{i}(\bm{d},\bm{\theta},\bm{\phi})$ satisfies an LN distribution with parameter $\sigma_{I}^{2}(d_{i})$ and $\{\eta_{0},\eta_{1},\cdots,\eta_{n}$ are independent variables, from (51) we can see that $\eta_{eq,n}(\bm{d},\bm{\theta},\bm{\phi})$ also satisfies an LN distribution with parameter $\sigma_{eq,n}^{2}(\bm{d},\bm{\theta},\bm{\phi})=\sum_{i=0}^{n}\sigma_{I}^{2}(d_{i})$. Therefore, the conditional PDF of $\eta_{eq,n}(\bm{d},\bm{\theta},\bm{\phi})$ is given by
$f_{H,n}(\eta_{eq,n}|\bm{d},\bm{\theta},\bm{\phi})$ $=\frac{1}{\eta_{eq,n}\sqrt{2\pi\sigma_{eq,n}^{2}(\bm{d},\bm{\theta},\bm{\phi})}}\exp\left(\!{-\frac{(\ln\eta_{eq,n}+0.5\sigma_{eq,n}^{2}(\bm{d},\bm{\theta},\bm{\phi}))^{2}}{2\sigma_{eq,n}^{2}(\bm{d},\bm{\theta},\bm{\phi})}}\!\right).$ (52)
Then the unconditional PDF for the equivalent turbulent fading coefficient $\eta_{eq,n}$ can be obtained by averaging $\{\bm{d},\bm{\theta},\bm{\phi}\}$ on $\mathcal{F}$ out, which is given by
$f_{H,n}(\eta_{eq,n})$ $=\frac{\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}\frac{I_{n}(\bm{d},\bm{\theta},\bm{\phi})}{\eta_{eq,n}\sqrt{2\pi\sum_{i=0}^{n}\sigma_{I}^{2}(d_{i})}}\exp\left({-\frac{(\ln\eta_{eq,n}+0.5\sum_{i=0}^{n}\sigma_{I}^{2}(d_{i}))^{2}}{2\sum_{i=0}^{n}\sigma_{I}^{2}(d_{i})}}\right)f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}}{\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}I_{n}(\bm{d},\bm{\theta},\bm{\phi})f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}}.$ (53)
Then we can also use the MCI approach to estimate $f_{H}(\eta_{eq,n})$ using the same sampling PDF given in (25). Using the same sampling points $\{\bm{s}_{n}^{1},\bm{s}_{n}^{2},\cdots,\bm{s}_{n}^{M}\}$, we can estimate the PDF of equivalent turbulent fading coefficient $f_{H}(\eta_{eq,n})$ as
$f_{H,n}(\eta_{eq,n})$ $=\frac{1}{Count}\sum_{m=1}^{M}\frac{I_{n}(\bm{s}_{n}^{m})}{\eta_{eq,n}\sqrt{2\pi\sum_{i=0}^{n}\sigma_{I}^{2}(d_{i}^{m})}}\exp\left({-\frac{(\ln\eta_{eq,n}+0.5\sum_{i=0}^{n}\sigma_{I}^{2}(d_{i}^{m}))^{2}}{2\sum_{i=0}^{n}\sigma_{I}^{2}(d_{i}^{m})}}\right).$ (54)
@@
$f_{H}(\eta_{eq})$ $=\frac{P_{tot}}{P_{1}}f_{H,1}\left(\frac{P_{tot}}{P_{1}}\eta_{eq}\right)*\frac{P_{tot}}{P_{2}}f_{H,2}\left(\frac{P_{tot}}{P_{2}}\eta_{eq}\right)*\cdots*\frac{P_{tot}}{P_{N}}f_{H,N}\left(\frac{P_{tot}}{P_{N}}\eta_{eq}\right),$ (55)
@@
The pseudocode of the MCI approach is similar to the one we presented in [17]. The only difference is that we need to add the estimation of turbulent fluctuation and distribution in each MCI process according to the equations given in (47) and (54), respectively.
V Numerical Results and Discussions
In this section, we present some numerical results to verify our theoretical analysis and discuss some interesting findings. Unless otherwise specified, the simulation parameters we used in this section are listed in Table I. Without loss of generality, we set $C_{n}^{2}=10^{-17}$ $\text{m}^{-2/3}$, $C_{n}^{2}=10^{-15}$ $\text{m}^{-2/3}$, and $C_{n}^{2}=10^{-13}$ $\text{m}^{-2/3}$ for weak, moderate, and strong turbulent conditions, respectively.

### v3

@@ IV-B Estimating the Turbulent Fluctuation
where we have used the fact that all sampling points $\{\bm{s}_{n}^{1},\bm{s}_{n}^{2},\cdots,\bm{s}_{n}^{M}\}$ are independent from each other and $\text{Var}[O_{n}(\bm{s}_{n}^{1})]=\text{Var}[O_{n}(\bm{s}_{n}^{2})]\cdots=\text{Var}[O_{n}(\bm{s}_{n}^{M})]=\text{Var}[O_{n}(\bm{d},\bm{\theta},\bm{\phi})]$ is the variance of the objective function.
From eq. (31), we can see that the variance $\text{Var}[P_{tur,n}]$ depends on both $\text{Var}[O_{n}(\bm{d},\bm{\theta},\bm{\phi})]$ and $M$. Besides, we have $\text{Var}[P_{tur,n}]\to 0$ when the number of sampling points $M\to\infty$, which means the calculated variance can be arbitrarily small as long as $M$ is large enough. This is because a larger number of sampling points $M$ in the Monte-Carlo process will reduce the fluctuation caused by randomness, leading to more stable and precise results. Therefore, the attempt to calculate the turbulent fluctuation by estimating $\text{Var}[P_{tur,n}]$ becomes unreliable due to the fact that $\text{Var}[P_{tur,n}]$ depends on $M$.
Here we present an improved method based on an MCI approach for estimating the turbulent fluctuation, which is more interpretable compared with the existing MCS approach and is independent of $M$ when $M$ is large.
We first substitute (14) into (16) and obtain the instantaneous receiving probability as
$P_{tur,n}(\bm{\eta},\bm{d},\bm{\theta},\bm{\phi})$ (32)
[…]
@@
$P_{tur,n}(\bm{\eta})$ $=\!\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}P_{n}(\bm{d},\bm{\theta},\bm{\phi})\!\prod_{i=0}^{n}\eta_{i}\!f(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}$ (33)
$=\!{\prod_{i=0}^{n}\eta_{i}}P_{n}.$
Then we can define an equivalent turbulent fading coefficient as
$\eta_{eq,n}\triangleq\frac{P_{tur,n}(\bm{\eta})}{P_{n}}$ (34)
on the average receiving power $P_{n}$ such that the instantaneous receiving power $P_{tur,n}(\bm{\eta})=\eta_{eq,n}P_{n}$. Obviously, we have
@@
$\sigma_{tur,n}^{2}$ $\triangleq\text{Var}[\eta_{eq,n}]=\int_{\bm{\eta}}(\eta_{eq,n}-1)^{2}f(\bm{\eta})\mathrm{d}\bm{\eta},$ (36)
where $f(\bm{\eta})$ is the joint PDF of $\{\eta_{0},\eta_{1},\cdots,\eta_{n}\}$. From (32) we can see that those propagating paths with $\{\bm{d},\bm{\theta},\bm{\phi}\}\notin\mathcal{F}$ contribute zero to the turbulent fluctuation effect. Therefore, we should restrict the variance on the set $\mathcal{F}$. Then the joint PDF $f(\bm{\eta})$ can be obtained by averaging $\{\bm{d},\bm{\theta},\bm{\phi}\}$ out of the conditional PDF $f(\bm{\eta}|\bm{d},\bm{\theta},\bm{\phi})$ in (18) on the set $\mathcal{F}$, i.e.,
$f_{H}(\bm{\eta})$ $\!=\!\int_{\mathcal{F}}f(\bm{d},\bm{\theta},\bm{\phi}|\{\bm{d},\bm{\theta},\bm{\phi}\}\in\mathcal{F})$ (37)
$\quad\quad\times f(\bm{\eta}|\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi},$
where $f(\bm{d},\bm{\theta},\bm{\phi}|\{\bm{d},\bm{\theta},\bm{\phi}\}\in\mathcal{F})$ is the joint PDF of $\{\bm{d},\bm{\theta},\bm{\phi}\}$ on the set $\mathcal{F}$; and we have
$f(\bm{d},\bm{\theta},\bm{\phi}|\{\bm{d},\bm{\theta},\bm{\phi}\}\in\mathcal{F})$ $=\frac{f(\bm{d},\bm{\theta},\bm{\phi})}{\text{Pr}[\{\bm{d},\bm{\theta},\bm{\phi}\}\in\mathcal{F}]},$ (38)
@@
$\text{Pr}[\{\bm{d},\bm{\theta},\bm{\phi}\}\in\mathcal{F}]=\int_{\mathcal{F}}f(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}.$ (39)
@@
$f_{H}(\bm{\eta})$ $=\frac{\int_{\mathcal{F}}\prod_{i=0}^{n}f(\eta_{i};d_{i})f(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}}{\int_{\mathcal{F}}f(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}}$ (40)
$=\frac{\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}\prod_{i=0}^{n}f(\eta_{i};d_{i})I_{n}f(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}}{\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}I_{n}f(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}}.$
@@
$\sigma_{tur,n}^{2}$ $\!=\!\frac{\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}(\prod_{i=0}^{n}M_{2}(d_{i})-1)I_{n}f(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}}{\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}I_{n}f(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}},$ (41)
@@
$\sigma_{tur,n}^{2}$ $=\frac{\frac{1}{M}\sum_{m=1}^{M}\left[\prod_{i=0}^{n}M_{2}(d_{i}^{m})-1\right]I_{n}(\bm{s}_{n}^{m})}{\frac{1}{M}\sum_{m=0}^{M}I_{n}(\bm{s}_{n}^{m})}$ (42)
$=\frac{1}{Count}\sum_{m=1}^{M}\left[\prod_{i=0}^{n}M_{2}(d_{i}^{m})-1\right]I_{n}(\bm{s}_{n}^{m}),$
@@
Similar to the $n$th scattering order, the total turbulent fluctuation effect on the total receiving power of $N$ scattering orders can be characterized by defining an equivalent turbulent fading coefficient as
$\eta_{eq}\triangleq\frac{\sum_{n=1}^{N}P_{tur,n}(\bm{\eta})}{\sum_{n=1}^{N}P_{n}},$ (43)
which allows us to treat the complex multiple-scattering channel as a single equivalent link with a unified fading parameter and simplify the characterization of turbulent fluctuation effects. Using the relations $P_{tur,n}(\bm{\eta})=\eta_{eq,n}P_{n}$ and $P_{tot}=\sum_{n=1}^{N}P_{n}$, we can obtain
@@
IV-C Estimating the Distribution of Turbulent Fading Coefficient
Now we further derive the distributions of the equivalent turbulent fading coefficients $\eta_{eq,n}$ and $\eta_{eq}$. For a given propagating path $\{\bm{d},\bm{\theta},\bm{\phi}\}$, the equivalent turbulent fading coefficient $\eta_{eq,n}$ can be obtained from (35) as
@@ IV-C Estimating the Distribution of Turbulent Fading Coefficient
When $\eta_{i}(\bm{d},\bm{\theta},\bm{\phi})$ satisfies an LN distribution with parameter $\sigma_{I}^{2}(d_{i})$ and $\eta_{0},\eta_{1},\cdots,\eta_{n}$ are independent variables, from (47) we can see that $\eta_{eq,n}(\bm{d},\bm{\theta},\bm{\phi})$ also satisfies an LN distribution with parameter $\sigma_{eq,n}^{2}(\bm{d},\bm{\theta},\bm{\phi})=\sum_{i=0}^{n}\sigma_{I}^{2}(d_{i})$. Therefore, the conditional PDF of $\eta_{eq,n}(\bm{d},\bm{\theta},\bm{\phi})$ is given by
$f_{H,n}(\eta_{eq,n}|\bm{d},\bm{\theta},\bm{\phi})$ $=\frac{\exp\left(\!{-\frac{(\ln\eta_{eq,n}+0.5\sigma_{eq,n}^{2}(\bm{d},\bm{\theta},\bm{\phi}))^{2}}{2\sigma_{eq,n}^{2}(\bm{d},\bm{\theta},\bm{\phi})}}\!\right)}{\eta_{eq,n}\sqrt{2\pi\sigma_{eq,n}^{2}(\bm{d},\bm{\theta},\bm{\phi})}}.$ (48)
Then the unconditional PDF for the equivalent turbulent fading coefficient $\eta_{eq,n}$ can be obtained by averaging $\{\bm{d},\bm{\theta},\bm{\phi}\}$ on $\mathcal{F}$ out. Specifically, we have
$f_{H,n}(\eta_{eq,n})$ $=\int_{\mathcal{F}}f(\bm{d},\bm{\theta},\bm{\phi}|\{\bm{d},\bm{\theta},\bm{\phi}\}\in\mathcal{F})$ (49)
$\quad\quad\times f_{H,n}(\eta_{eq,n}|\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi},$
where $f(\bm{d},\bm{\theta},\bm{\phi}|\{\bm{d},\bm{\theta},\bm{\phi}\}\in\mathcal{F})$ is given in (38). By submitting eqs. (38) and (39) into (49), we can obtain
$f_{H,n}(\eta_{eq,n})$ (50)
$\quad=\frac{\int_{\mathcal{F}}f(\bm{d},\bm{\theta},\bm{\phi})f_{H,n}(\eta_{eq,n}|\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}}{\int_{\mathcal{F}}f(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}}$
$\quad=\frac{\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}f_{H,n}(\eta_{eq,n}|\bm{d},\bm{\theta},\bm{\phi})I_{n}f(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}}{\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}I_{n}f(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}},$
where we have used the equality $\int_{\mathcal{F}}g(\bm{d},\bm{\theta},\bm{\phi})f(\bm{d},\bm{\theta},\bm{\phi})$ $\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}=\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}I_{n}g(\bm{d},\bm{\theta},\bm{\phi})f(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}$ for any function $g(\bm{d},\bm{\theta},\bm{\phi})$. This is because $I_{n}=0$ if $(\bm{d},\bm{\theta},\bm{\phi})$ is not in $\mathcal{F}$. Therefore, by substituting (48) into (50), we can finally obtain the unconditional PDF for the equivalent turbulent fading coefficient $\eta_{eq,n}$, which is given in (51) on the top of next page.
$f_{H,n}(\eta_{eq,n})$ $=\frac{\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}\frac{I_{n}}{\eta_{eq,n}\sqrt{2\pi\sum_{i=0}^{n}\sigma_{I}^{2}(d_{i})}}\exp\left({-\frac{(\ln\eta_{eq,n}+0.5\sum_{i=0}^{n}\sigma_{I}^{2}(d_{i}))^{2}}{2\sum_{i=0}^{n}\sigma_{I}^{2}(d_{i})}}\right)f(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}}{\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}I_{n}f(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}}.$ (51)
Then we can also adopt the MCI approach to estimate $f_{H,n}(\eta_{eq,n})$ by using the same sampling PDF given in (23). Using the same sampling points $\{\bm{s}_{n}^{1},\bm{s}_{n}^{2},\cdots,\bm{s}_{n}^{M}\}$, we can estimate the PDF of equivalent turbulent fading coefficient $f_{n}(\eta_{eq,n})$ as
$f_{H,n}(\eta_{eq,n})$ $=\frac{1}{Count}\sum_{m=1}^{M}\frac{I_{n}(\bm{s}_{n}^{m})}{\eta_{eq,n}\sqrt{2\pi\sum_{i=0}^{n}\sigma_{I}^{2}(d_{i}^{m})}}$ (52)
$\times\exp\left({-\frac{(\ln\eta_{eq,n}+0.5\sum_{i=0}^{n}\sigma_{I}^{2}(d_{i}^{m}))^{2}}{2\sum_{i=0}^{n}\sigma_{I}^{2}(d_{i}^{m})}}\right).$
@@
$f_{H}(\eta_{eq})$ $=\frac{P_{tot}}{P_{1}}f_{H,1}\left(\frac{P_{tot}}{P_{1}}\eta_{eq}\right)*\frac{P_{tot}}{P_{2}}f_{H,2}\left(\frac{P_{tot}}{P_{2}}\eta_{eq}\right)$ (53)
$\quad*\cdots*\frac{P_{tot}}{P_{N}}f_{H,N}\left(\frac{P_{tot}}{P_{N}}\eta_{eq}\right),$
@@
The pseudocode of the MCI approach is similar to the one we presented in [yuan2019monte]. The only difference is that we need to add the estimation of turbulent fluctuation and distribution in each MCI process according to the equations given in (42) and (52), respectively.
We remark that though our derivation here is based on LN distribution for weak turbulence, we can also extend this derivation to moderate and strong turbulence cases. However, in moderate and strong turbulence, a large $C_{n}^{2}$ can result in a large Rytov variance $\sigma_{r}^{2}$ and thus a large turbulent variance $\text{Var}[\eta]\gg 1$ if LN distribution is adopted, which does not match practical moderate and strong turbulence with $\text{Var}[\eta]$ normally less than 2. An alternative way to extend our method to moderate and strong turbulence is using the Gamma-Gamma fading to calculate the turbulent variance $\text{Var}[\eta]$ since it never can never exceed 2; then we use this turbulent variance $\text{Var}[\eta]$ to calculate the parameter $\sigma_{I}^{2}$ in LN fading as $\sigma_{I}^{2}=\ln(\text{Var}[\eta]+1)$. Through this way we can estimate the distribution of turbulent fading coefficient for all turbulent conditions.
V Simulation and Experimental Results
In this section, we present some simulation and experimental results to verify our theoretical analysis. Unless otherwise specified, the simulation parameters we used in this section are listed in Table I. Without loss of generality, we set $C_{n}^{2}=10^{-17}$ $\text{m}^{-2/3}$, $C_{n}^{2}=10^{-15}$ $\text{m}^{-2/3}$, and $C_{n}^{2}=10^{-13}$ $\text{m}^{-2/3}$ for weak, moderate, and strong turbulent conditions, respectively.